# Parabolic bundles and the intersection cohomology of moduli spaces of vector bundles on curves

### v1

@@ 5.2. The E-polynomial of non-abelian fibers
To show that the injection (43) is an isomorphism, it is enough to prove the following.
@@ 5.2.1. The right hand side
Without loss of generality, we assume that $i=k$. Recall that in the proof of Theorem 4.12 we identified $\mathcal{F}_{k}$ with the projective bundle over the ”smaller” fiber of $\pi$ (see (37)). Thus using induction we can conclude that $H^{\bullet}(\mathcal{F}_{k})$ is generated by the $k$ classes $x_{1},\dots,x_{k}\in H^{2}(\mathcal{F}_{k})$. It follows from Lemma 4.15 that these classes are well-defined in the cohomology ring $H^{\bullet}(F_{V})$, hence the restriction map $H^{\bullet}(F_{V})\rightarrow H^{\bullet}(\mathcal{F}_{k})$ is surjective. ∎
@@
  • with some abuse of notation, for any $J\subset\underline{k}$, we denote by $\kappa_{j}:H^{\bullet}(\mathcal{F}_{J\setminus\{j\}})\longrightarrow H^{\bullet}(\mathcal{F}_{J})$ the surjection from Lemma 5.7 if $j\in J$, and zero otherwise.
@@
acting on the direct sum of vector spaces $H^{\bullet}_{J}=e_{J}\otimes H^{\bullet}(\mathcal{F}_{J})$.
@@
Indeed, recall that the spaces $\mathcal{F}_{J}$ are bundles of products of projective spaces over spaces of the form $F_{V^{\prime}}$. Since these latter spaces have cohomology with pure Hodge structure (see Lemma 4.11), so do the spaces $\mathcal{F}_{J}$. This immediately implies the exactness of (50). ∎
@@
where we denote by $\mathcal{U}_{k-l}$ the normalized (cf. def. 3.9) universal bundle over $F_{(k-l)W}\times C$, by $\xi$ and $\zeta$ the projections along $C$ and $F_{(k-l)W}$, respectively, and treat $W$ as a parabolic bundle with trivial parabolic structure. For the last term in (51) we have $H_{\underline{k}}^{\Sigma_{k}}\cong\wedge^{k}H^{\bullet}(\prod_{i=1}^{k}\mathbb{P}(W_{p}))$.
@@
First, we observe that $\oplus_{|J|=l}H^{\bullet}_{J}$, as a representation of $\Sigma_{k}$, is induced from the representation $H^{\bullet}_{\{1,2,\dots,l\}}$ of the stabilizer subgroup $\Sigma_{l}\times\Sigma_{k-l}$, and thus by Frobenius reciprocity, we have
@@
$\bigotimes_{j=1}^{l}H^{\bullet}(\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}V_{j})))\otimes H^{\bullet}(F_{(k-l)W}),$
@@
(52) $\wedge^{l}H^{\bullet}(\mathbb{P}R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}W))\otimes H^{\bullet}(F_{(k-l)W}).$
@@ 5.2.2. The left hand side
=$\mathcal{E}\overset{\mathrm{def}}{=}\{0\to W\to W^{\prime\prime}\to W^{\prime}\to 0|\;W^{\prime}\in\pi^{-1}(W),\;W^{\prime\prime}\in\mathcal{P}_{0}(2s)\}$
@@
As opposed to the abelian case, some of parabolic extensions $(W^{\prime\prime},l)$ obtained in such way, are isomorphic, however. Indeed, consider the locus $\mathcal{G}\subset\mathcal{E}$ of those parabolic bundles $W^{\prime\prime}$ which, as vector bundles, are isomorphic to $W\oplus W$. The relevant parabolic structure on $W^{\prime\prime}\in\mathcal{G}$ is a choice of a line $l$ in $W_{p}$, and then an appropriate lift of this line to $W_{p}\oplus W_{p}$. Thus it is easy to see that the locus $\mathcal{G}\subset\mathcal{E}$ forms the space of (1,2)-flags of $W_{p}$.
@@
The final picture of the fiber $F_{W\oplus W}$ is the following: over a base $\mathbb{P}^{s-1}$, there is a projective bundle with fibers $\mathbb{P}^{s^{2}(g-1)+s-1}$, which contains a subbundle $\mathcal{G}$, which forms a (1,2)-flag variety of the vector space $W_{p}$. The projection $\mathcal{G}\to\mathrm{Gr}_{2}(W_{p})$ is a $\mathbb{P}^{1}$-fibration, and to obtain the fiber $F_{W\oplus W}$, we need to collapse this fibration to the base.
@@ 6. Combinatorics of subtraction
Moving to the next step of our plan (see §3.6), in this section we begin the study of the stalk of the local system over the abelian stratum $M_{\rho}^{\mathrm{ab}}$ contributing to the Decomposition Theorem.
@@
  • $L_{\rho}=(\mathcal{L}_{\rho})_{V}$ for the stalk of the local system supported at $V$; we thus have $L_{\rho}\subset H^{\bullet}(F_{\rho})$.
Our goal in this section is to calculate the Hilbert function of the graded vector space $L_{\rho}\subset H^{\bullet}(F_{\rho})$, using equality (1.7.B) of the Decomposition Theorem. Applied to our situation the formula has the form
@@
For $\rho=[r_{1},\ldots,r_{k}]\vdash r$, the graded vector space $L_{\rho}\subset H^{\bullet}(F_{\rho})$ is the stalk of the local system $\mathcal{L}_{\rho}$ supported on the stratum $M_{\rho}^{\mathrm{ab}}$. Then the Hilbert function $\mathrm{Hilb}_{t}(L_{\rho})=\sum_{i}\dim L^{i}_{\rho}\cdot t^{i}$ is given by
@@ 6.1. The subtraction algorithm
Given a polystable bundle $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ and $\underline{\lambda}\vdash\underline{k}$, then we can define the bundles
@@
The following technical statements are straightforward and their proof will be omitted.
Lemma 6.3.
Suppose $\mu=\underline{\lambda}\circ\rho$, and $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$. Then $M_{\rho}\subset\overline{M}_{\mu}$ and
  1. (1)
@@
  3. (3)
    The local system $\mathcal{L}_{\mu}$ is trivial on $M_{\mu}$ in a neighborhood of $V$.
These statements imply that
$IH^{\bullet}(\mathcal{C}_{\mu\rho},\mathcal{L}_{\mu})=\sum_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}L_{\mu}\otimes\bigotimes_{j=1}^{n}IH^{\bullet}(\mathcal{C}_{\rho|\underline{\lambda}_{j}}).$
@@ 6.2. Acyclic subgraphs and intersection cohomology
Finally, by generating function of a set of graphs $\mathcal{A}$ we will always mean counting the graphs weighted by the number of their edges: usually, of the form $t^{N}\cdot\sum_{\Gamma\in\mathcal{A}}t^{e(\Gamma)}$ for some integer $N$.
@@
  • We denote by $\mathbb{F}^{\circ}_{\rho}\subset\mathbb{F}_{\rho}$ the subgraph obtained from $\mathbb{F}_{\rho}$ by removing the vertex labelled by 00 and all edges attached to this vertex.
@@ 6.3. Poincaré polynomials of the fibers in terms of graphs
  • the graph $G\setminus J$ obtained from $G$ by deleting all vertices labelled by $j\in J$ and all edges attached to these vertices is an acyclic subgraph of $\mathbb{F}_{\rho}$ rooted at zero vertex, with vertex set $\overline{J}\cup\{0\}$;
@@
We are now ready to present the proof of Theorem 6.1: Substituting $g(\rho;t^{2})=IP_{t}(\mathcal{C}_{\rho},t)$ and $f(\rho;t^{2})=P_{t}(F_{\rho})$ (cf. Theorems 6.6 and 6.7) into (61), we obtain a recursion
@@
The key point is an observation of [31, Theorem 6.3]: the function $\hat{g}(\Gamma,t^{2})$ coincides with the Hilbert function of the quotient algebra
@@
7. Calculation of the local systems
@@ 7.1. Refined intersection forms
which is exact in the middle, i.e. $\operatorname{im}i_{*}\cong\ker j^{*}$. Now we can conclude the following:
@@ 7.3. Interlude on excess intersection formulas
[13, §13.3] Let $S$ and $T$ be smooth projective subvarieties of a smooth variety $X$ with $\dim S+\dim T=\dim X$. Then
@@
Now we will evaluate the element $w_{\sigma,\tau}$. Using Propositions 4.9 and 4.19 we obtain that
@@ 8.1. Quotients and local systems
Our first step is to calculate the monodromy of the local system $\mathcal{L}_{\rho}$. For a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$, we introduce the subgroup $\mathrm{Aut}(\rho)\subset\Sigma_{k}$:
@@
and the subvariety
$\widetilde{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},V_{2},\dots,V_{k})|\,V_{i}\in\mathcal{M}_{0}(r_{i})^{\mathrm{st}},V_{i}\not\cong V_{j}\text{ if }i\neq j\}\subset\\ \mathcal{M}_{0}(r_{1})^{\mathrm{st}}\times\mathcal{M}_{0}(r_{2})^{\mathrm{st}}\times\dots\times\mathcal{M}_{0}(r_{k})^{\mathrm{st}},$
such that $M_{\rho}^{\mathrm{ab}}=\widetilde{M}_{\rho}^{\mathrm{ab}}/\mathrm{Aut}(\rho)$.
@@
$\mathcal{L}_{\rho}=\widetilde{M}_{\rho}^{\mathrm{ab}}\times_{\mathrm{Aut}(\rho)}L_{\rho}.$
Now we pass to the intermediate extensions. We will need the following basic facts from the theory (see [14, §4.2] and [20, Lemma 2.12]).
@@
$\mathbb{S}^{n}[-q]\cong\begin{cases}\mathbb{Q}[-nq]\text{ if }n\text{ is even},\\ \mathbb{S}[-nq]\text{ if }n\text{ is odd}.\end{cases}$
@@ 8.2. The plethystic exponential and the final result
Next, combining Proposition 8.2 and Corollary 8.6, we arrive at our main result:
@@ 8.3. Variants and refinements
Finally, we show that our formulas can be generalized to describe the Hodge structuure on $IH^{\bullet}(\mathcal{M}_{0}(r))$. For a smooth, projective variety $Y$, denote the generator of its Hodge numbers by
@@
Because of purity, setting $u=v=t$, we recover the E-polynomial (cf. 3.18): $E_{t}(Y)=h_{t,t}(Y)$

### v2

@@ 6.3. The E-polynomial of non-abelian fibers
To show that the injection (44) is an isomorphism, given the purity of the cohomology of the fibers, it is enough to prove the following. Recall the definition of weight polynomial in §4.2.
@@ 6.3.1. The right hand side
Without loss of generality, we assume that $i=k$. Recall that in the proof of Theorem 5.13 we identified $\mathcal{F}_{k}$ with the projective bundle over the ”smaller” fiber of $\pi$ (see (38)). Thus, using induction, we can conclude that $H^{\bullet}(\mathcal{F}_{k})$ is generated by the $k$ classes $x_{1},\dots,x_{k}\in H^{2}(\mathcal{F}_{k})$. It follows from Lemma 5.17 that these classes are well-defined in the cohomology ring $H^{\bullet}(F_{V})$, hence the restriction map $H^{\bullet}(F_{V})\rightarrow H^{\bullet}(\mathcal{F}_{k})$ is surjective. ∎
@@
  • with some abuse of notation, for any $J\subset\underline{k}$, we let $\kappa_{j}:H^{\bullet}(\mathcal{F}_{J\setminus\{j\}})\longrightarrow H^{\bullet}(\mathcal{F}_{J})$ be the surjection from Lemma 6.7 if $j\in J$, and zero otherwise.
@@
acting on the direct sum of vector spaces $H^{\bullet}_{J}=e_{J}\otimes H^{\bullet}(\mathcal{F}_{J})$. This operator extends ${\bar{Q}}$ defined in (48).
@@
Indeed, recall that the spaces $\mathcal{F}_{J}$ are bundles of products of projective spaces over spaces of the form $F_{V^{\prime}}$. Since these latter spaces have pure cohomology (see Proposition 5.10), so do the spaces $\mathcal{F}_{J}$. This immediately implies the exactness of (52). ∎
@@
where we denote by $\mathcal{U}_{k-l}$ the normalized (cf. def. 3.12) universal bundle over $F_{(k-l)W}\times C$, by $\zeta$ and $\xi$ the projections along $C$ and $F_{(k-l)W}$, respectively, and treat $W$ as a parabolic bundle with trivial parabolic structure. For the last term in (53) we have $H_{\underline{k}}^{\Sigma_{k}}\cong\wedge^{k}H^{\bullet}(\prod_{i=1}^{k}\mathbb{P}(W_{p}))$.
@@
First, we observe that, as a representation of $\Sigma_{k}$, $\oplus_{|J|=l}H^{\bullet}_{J}$ is induced from the representation $H^{\bullet}_{\{1,2,\dots,l\}}$ of the stabilizer subgroup $\Sigma_{l}\times\Sigma_{k-l}$, and thus by Frobenius reciprocity, we have
@@
$\left[H^{\bullet}_{\{1,2,\dots,l\}}\right]^{\Sigma_{k-l}}\cong\bigotimes_{j=1}^{l}H^{\bullet}(\mathbb{P}(R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}V_{j})))\otimes H^{\bullet}(F_{(k-l)W}),$
@@
(54) $\left[H^{\bullet}_{\{1,2,\dots,l\}}\right]^{\Sigma_{l}\times\Sigma_{k-l}}\cong\wedge^{l}H^{\bullet}(\mathbb{P}R^{1}\xi_{*}\mathrm{ParHom}(\mathcal{U}_{k-l},\zeta^{*}W))\otimes H^{\bullet}(F_{(k-l)W}).$
@@ 6.3.2. The left hand side
=$\mathcal{E}{=}\{0\to W\to W^{\prime\prime}\to W^{\prime}\to 0|\;W^{\prime}\in\pi^{-1}(W),\;W^{\prime\prime}\in\mathcal{P}_{0}(2s)\}$
@@
As opposed to the abelian case, some of parabolic extensions $(W^{\prime\prime},l)$ obtained in such way, are isomorphic, however. Indeed, consider the locus $\mathcal{G}\subset\mathcal{E}$ of those parabolic bundles $W^{\prime\prime}$ which, as vector bundles, are isomorphic to $W\oplus W$. The relevant parabolic structure on $W^{\prime\prime}\in\mathcal{G}$ is a choice of a line $l$ in $W_{p}$, and then an appropriate lift of this line to $W_{p}\oplus W_{p}$. Thus the locus $\mathcal{G}\subset\mathcal{E}$ forms the space of (1,2)-flags of $W_{p}$.
@@
The final picture of the fiber $F_{W\oplus W}$ is the following: over a base $\mathbb{P}^{s-1}$, there is a projective bundle with fibers $\mathbb{P}^{s^{2}(g-1)+s-1}$, which contains a subbundle $\mathcal{G}$, which forms a (1,2)-flag variety of the vector space $W_{p}$. The projection $\mathcal{G}\to\mathrm{Gr}_{2}(W_{p})$ is a $\mathbb{P}^{1}$-fibration, and to obtain the fiber $F_{W\oplus W}$, we need to collapse this fibration to the base. ∎
@@ 7. Combinatorics of subtraction
Moving to the next step of our plan (see §3.5), in this Section, we begin the study of the stalk of the local system contributing to the Decomposition Theorem over an abelian point.
@@
  • $L_{\rho}=(\mathcal{L}_{\rho})_{V}$ for the stalk of the local system supported at $V$, which forms an ideal $L_{\rho}\subset H^{\bullet}(F_{\rho})$.
Our goal is to calculate the Hilbert function of the graded vector space $L_{\rho}\subset H^{\bullet}(F_{\rho})$, using equality (1.7.B) of the Decomposition Theorem. Applied to our situation the formula has the form
@@
For $\rho=[r_{1},\ldots,r_{k}]\vdash r$. The Hilbert function of $L_{\rho}$ is given by
@@ 7.1. The subtraction algorithm
Given a bundle $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ and $\underline{\lambda}\vdash\underline{k}$, then we can define the bundles
@@
Suppose $\mu=\underline{\lambda}\circ\rho$, and $V=V_{1}\oplus\ldots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$. Then $M_{\rho}\subset\overline{M}_{\mu}$ and it is straightforward to see that
@@
Since the local system $\mathcal{L}_{\mu}$ is trivial on $M_{\mu}$ in a neighborhood of $V$, we can conclude the following technical statement.
Lemma 7.3.
$IH^{\bullet}(\mathcal{C}_{\mu\rho},\mathcal{L}_{\mu})=\bigoplus_{\underline{\lambda}:\,\mu=\underline{\lambda}\circ\rho}L_{\mu}\otimes\bigotimes_{j=1}^{n}IH^{\bullet}(\mathcal{C}_{\rho|\underline{\lambda}_{j}}).$
@@ 7.2. Acyclic subgraphs and intersection cohomology
Finally, by generating function of a set of graphs $\mathcal{A}$ we will always mean the one counting the graphs weighted by the number of their edges: usually, of the form $t^{N}\cdot\sum_{\Gamma\in\mathcal{A}}t^{e(\Gamma)}$ for some integer $N$.
@@
  • We denote by $\mathbb{F}^{\circ}_{\rho}\subset\mathbb{F}_{\rho}$ the subgraph obtained from $\mathbb{F}_{\rho}$ by removing the vertex labeled by 00 and all edges attached to this vertex.
@@ 7.3. Poincaré polynomials of the fibers in terms of graphs
  • the graph $G\setminus J$ obtained from $G$ by deleting all vertices labeled by $j\in J$ and all edges attached to these vertices is an acyclic subgraph of $\mathbb{F}_{\rho}$ rooted at zero vertex, with vertex set $\overline{J}\cup\{0\}$;
@@
We are now ready to present the proof of Theorem 7.1: Substituting $g(\rho;t^{2})=IP_{t}(\mathcal{C}_{\rho},t)$ and $f(\rho;t^{2})=P_{t}(F_{\rho})$ (cf. Theorems 7.6 and 7.7) into (63), we obtain a recursion
@@
The key point is an observation of [36, Theorem 6.3]: the function $\hat{g}(\Gamma,t^{2})$ coincides with the Hilbert function of the quotient algebra
@@
8. Refined intersection forms and the calculation of the local systems
@@ 8.1. Refined intersection forms
which is exact in the middle, i.e. $\operatorname{im}i_{*}\cong\ker j^{*}$. Now we can conclude the following (see also, [7]):
@@ 8.3. Interlude on excess intersection formulas
[14, §13.3] Let $S$ and $T$ be smooth projective subvarieties of a smooth variety $X$ with $\dim S+\dim T=\dim X$. Then
@@
Now we will evaluate the element $w_{\sigma,\tau}$. Using Propositions 5.8 and 5.21 we obtain that
@@ 9.1. Quotients and local systems
Our first step is to calculate the monodromy of the local system $\mathcal{L}_{\rho}$. For a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$, we recall the definition of the subgroup $\mathrm{Aut}(\rho)\subset\Sigma_{k}$:
@@
and the variety
$\mathcal{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},V_{2},\dots,V_{k})|\,V_{i}\in\mathcal{M}_{0}(r_{i})^{\mathrm{st}},V_{i}\not\cong V_{j}\text{ if }i\neq j\}\subset\\ \mathcal{M}_{0}(r_{1})^{\mathrm{st}}\times\mathcal{M}_{0}(r_{2})^{\mathrm{st}}\times\dots\times\mathcal{M}_{0}(r_{k})^{\mathrm{st}},$
such that $M_{\rho}^{\mathrm{ab}}=\mathcal{M}_{\rho}^{\mathrm{ab}}/\mathrm{Aut}(\rho)$.
@@
$\mathcal{L}_{\rho}=\mathcal{M}_{\rho}^{\mathrm{ab}}\times_{\mathrm{Aut}(\rho)}L_{\rho}.$
Now we pass to the intermediate extensions. We will need the following basic facts from the theory (see [15, §4.2] and [22, Lemma 2.12]).
@@
$\mathbb{S}^{n}[-q]\cong\begin{cases}\mathbb{Q}[-nq]\quad\mbox{if $n$ is even},\\ \mathbb{S}[-nq]\quad\mbox{if $n$ is odd}.\end{cases}$
@@ 9.2. The plethystic exponential and the final result
Next, combining Theorem 9.2 and Corollary 9.6, we arrive at our main result:
@@ 9.3. Variants and refinements
Finally, we show that our formulas can be generalized to describe the Hodge structure on $IH^{\bullet}(\mathcal{M}_{0}(r))$. For a smooth, projective variety $Y$, denote the generator of its Hodge numbers by
@@
Because of purity, setting $u=v=t$, we recover the E-polynomial (cf. 4.4): $E_{t}(Y)=h_{t,t}(Y)$